### v1

# InternVideo2: Scaling Video Foundation Models for Multimodal Video Understanding

Abstract
We introduce InternVideo2, a new video foundation model (ViFM) that achieves the state-of-the-art performance in action recognition, video-text tasks, and video-centric dialogue. Our approach employs a progressive training paradigm that unifies the different self- or weakly-supervised learning frameworks of masked video token reconstruction, cross-modal contrastive learning, and next token prediction. Different training stages would guide our model to capture different levels of structure and semantic information through different pretext tasks. At the data level, we prioritize the spatiotemporal consistency by semantically segmenting videos and generating video-audio-speech captions. This improves the alignment between video and text. We scale both data and model size for our InternVideo2. Through extensive experiments, we validate our designs and demonstrate the state-of-the-art performance on over 60 video and audio tasks. Notably, our model outperforms others on various video-related captioning, dialogue, and long video understanding benchmarks, highlighting its ability to reason and comprehend long temporal contexts.
00footnotetext: *Equal contribution. Corresponding authors.
1 Introduction
The pursuit of transferable spatiotemporal representations is a critical research area within vision understanding, holding diverse applications across domains such as video searching Gabeur et al. (2020), game control Bruce et al. (2024), robotic learning Driess et al. (2023), self-driving Zablocki et al. (2022), and scientific studies Team et al. (2023). Recently, the advancement of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023a); Touvron et al. (2023a, b) and their multimodal variations (MLLMs) OpenAI (2023b); Gong et al. (2023); Liu et al. (2023); Zhu et al. (2023a) has had a profound impact on vision research and other disciplines, thanks to their learned world models OpenAI (2023c); Yang et al. (2024). Embedding videos effectively into these world models and harnessing their capabilities to enhance video embedding have emerged as pivotal tasks for advancing video understanding.
Previous research has identified several effective learning schemes for video representation, including reconstructing videos with masked inputs (MAE He et al. (2022) and VideoMAE Wang et al. (2023a)), aligning videos with languages Xu et al. (2021); Yan et al. (2022); Li et al. (2023a), and predicting the next token using videos Alayrac et al. (2022); Sun et al. (2023a); Li et al. (2023b). These approaches have been shown to be complementary and can be unified through a progressive training scheme. Notably, methods such as UMT Li et al. (2023a), InternVideo Wang et al. (2022), and VideoPrism Zhao et al. (2024) have utilized a two-stage training approach involving masked reconstruction and multimodal contrastive learning, leading to improved performance in downstream tasks. These studies have demonstrated the benefits of joint learning for enhancing spatiotemporal perception and video-language alignment. This paper extends upon this scheme by incorporating video-based next token prediction and scaling the entire training process, including models and data.
The proposed video foundation model InternVideo2, is built through a progressive learning scheme. The learning involves three stages aimed at improving the model’s spatiotemporal perception, aligning it with semantics from other modalities, and enhancing its world modeling through next token prediction Alayrac et al. (2022). In the initial stage of masked video token reconstruction, the model learns to reconstruct masked video tokens, allowing the video encoder to develop basic spatiotemporal perception. To estimate the masked tokens, vision encoders (InternViT Chen et al. (2023a) and VideoMAE-g Wang et al. (2023a)) trained differently are employed as proxies. In the next stage of multimodal learning, the architecture is expanded to include audio and text encoders. This not only improves the alignment between videos and text but also grants InternVideo2 the ability to handle video-audio tasks. By incorporating these additional modalities, the model’s understanding of videos is enriched and aligned with the semantics provided by audios. Finally, in the next-token prediction stage, a video-centric dialogue system and the corresponding instruction-finetuning dataset are leveraged to train InternVideo2. This transfer learning process allows the model to benefit from the knowledge acquired by LLM and others. By connecting InternVideo2 to LLMs, the video encoder is further updated through next-token prediction training, enhancing its ability to generate contextually appropriate next tokens.
In pretraining, we emphasize incorporating spatiotemporal consistency in the data. We utilize a large-scale multimodal video-centric dataset consisting of 412M data entries, which includes 2M videos, 50M video-text pairs, 60M video-audio-speech-text pairs, and 300M image-text pairs. In multimodal pretraining data, we segment videos into clips semantically, and we focus on recalibrating the descriptions with videos using three modalities: audio, video, and speech. To this end, we generate captions for these three modalities separately. Then individual captions are fused together to create a more comprehensive description. By incorporating audio, video, and speech modalities in the training process, we provide a richer and more detailed understanding of the videos. This holistic data approach to video-text alignment improves the model’s ability to comprehend and interpret the video accurately.
We evaluate InternVideo2 across a range of video-related tasks. These tasks span from basic spatiotemporal perception, such as action recognition, to more complex reasoning tasks, such as long video or procedure-aware question-answering (QA), as given in Fig. 1. The results (in Sec. 5) demonstrate that InternVideo2 achieves state-of-the-art performance on multiple tasks, able to analyzes and reasons sequences of actions. These highlight that InternVideo2 excels not only in video perception and video-language alignment but also in modeling the world. Its top performance on various benchmarks in the field of multimodal language models (MLLM) signifies its capability to effectively capture and understand video content. These empirical findings validate that InternVideo2 serves as a qualified video encoder for future exploration in video understanding.
In summary, our contributions to video understanding are as follows.
This paper introduces InternVideo2, a competitive video foundation model that leverages masked reconstruction, crossmodal contrastive learning, and next token prediction to make model perceptive, semantic, and capable of reasoning in video understanding. It achieves the state-of-the-art performance in 65 out of 74 video / audio tasks. Our model demonstrates superior performance in video-related dialogue and long video understanding, highlighting its potential in various world model studies and applications.
We provide an enhanced dataset to train InternVideo2. This includes the validation and incorporation of audio data during training, as well as the implementation of annotation techniques. These improvements result in significant enhancements in model generalization and performance.
2 Related Work
Video Understanding. Studies on learning video foundation models becomes increasingly crucial considering its wide applications Li and Wang (2020); Xu et al. (2021); Li et al. (2023a, 2022a); Hu et al. (2022); Dou et al. (2022); Shen et al. (2021); Yao et al. (2021); Sun et al. (2019); Zhu and Yang (2020); Wang et al. (2022); Chen et al. (2022a); Zellers et al. (2021, 2022); Zeng et al. (2023a, b); Chen et al. (2023b, c); He et al. (2023); Chen et al. (2024a); Zhao et al. (2024); Wang et al. (2023b); Yan et al. (2022); Feichtenhofer et al. (2022); Wang et al. (2023c); Miech et al. (2020); Tong et al. (2022); Wang et al. (2023a); Li et al. (2023c). Typical methods in building video foundation models (ViFM) include video-text contrastive learning Xu et al. (2021); Wang et al. (2023c, 2022), masked video modeling Tong et al. (2022); Wang et al. (2023a, 2022); Fu et al. (2021), and next token prediction Alayrac et al. (2022); Sun et al. (2023a, b). Notably, All-in-one Wang et al. (2023c) utilizes a single backbone with unified multiple pretraining objectives. On the other hand, UMT Li et al. (2023a) combines masked modeling with video-text contrastive learning, demonstrating strong performance in both action recognition and video-language tasks. Another approach is mPLUG-2 Xu et al. (2023), which introduces a new design for modulating different modalities. It shares a common module across modalities to enhance relations while incorporating modality-specific modules for discrimination. In addition to video-text pretraining, researchers have also explored the use of audio information in videos to improve performance. MERLOT Reserve Zellers et al. (2022) learns video representations using a large-scale dataset of video-speech-transcript pairs. VALOR Chen et al. (2023c) employs independent encoders for video, audio, and text and trains a joint visual-audio-text representation. VAST Chen et al. (2024a) constructs an audio-visual-speech dataset and develops a multimodal backbone that excels in video-audio-related tasks. VideoPrism Zhao et al. (2024) combines video-text contrastive learning and video token reconstruction on a combination of public and proprietary videos, achieving leading results across various video tasks.
Multimodal Large Language Models. With advances in large language models (LLMs) Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020); Wei et al. (2021); Chowdhery et al. (2022); Sun et al. (2023a, b), their multimodal versions (MLLMs) gain a lot heat considering its open-world visual task execution by simple instructions. Seminal works like Flamingo Alayrac et al. (2022) shows outstanding zero/few-shots performances over a range of multimodal tasks Goyal et al. (2017a); Plummer et al. (2015); Xu et al. (2016); Marino et al. (2019). Public MLLMs Zhu et al. (2023b); Liu et al. (2023); Gong et al. (2023) e.g. LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023) give visual instruction-tuning data for dialogue experience. Some video-related MLLM (VideoLLM) VideoChatGPT Maaz et al. (2023a) and Valley Luo et al. (2023) utilize ChatGPT to generate video instruction-tuning data, aiming to enhance instruction-following capabilities for real-world video comprehension.
3 Method
We learn InternVideo2 in three stages, illustrated in Fig. 2. The stages include spatiotemporal token reconstruction, video-audio-speech-language contrastive learning, and joint training of a large language model (LLM).
Video Encoder.
The video encoder used in InternVideo2 follows the Vision Transformer (ViT) Dosovitskiy et al. (2020) and includes additional projection layers for distillation. Inspired by previous works Chen et al. (2023a); Yu et al. (2022), we introduce attention pooling to the ViT. For input videos, we sparsely sample 8 frames and perform a 1414 () spatial downsampling. These spatiotemporal tokens are then concatenated with a class token and combined with 3D position embeddings. The hyperparameters used for ViT-1B and -6B are the same as CoCa-1B Yu et al. (2022) and InternViT-6B Chen et al. (2023a), respectively. The details of ViT-6B architecture are given in Supp. Tab. 21.
3.1 Stage1: Reconstructing Masked Video Tokens
We exploit two expert models to guide the video encoder to conduct masked token-level reconstruction. Specifically, we adopt InternVL-6B Chen et al. (2023a) and VideoMAEv2-g Wang et al. (2023a) to transfer unmasked knowledge via simple projection layers. When training, we input the full videos into different teachers and masked out 80% of the tokens frame by frame, under the semantic guidance of the multimodal model InternVL and motion-aware model VideoMAEv2. We only align the related unmasked tokens, by minimizing their mean squared error (MSE). The learning objective of this stage is formed by reconstructing the remaining tokens as:
where , , and are our video encoder, InternViT-6B Chen et al. (2023a), and ViT-g of VideoMAEv2, respectively. stands for the token index and is the corresponding token extracted from InternVideo2 of input video . is the normalization factor. and balance the influence between the employed models.
In implementation, we align the last layers of InternVL, last layers of VideoMAE, and the final output token of InternVL. Different losses are simply summed for optimization. After pretraining, we drop those projection layers and only use the basic encoder. Compared with only using the multimodal model in UMT and VideoPrism, our strategy makes the vision encoder multimodal-friendly as well as enhances its temporal sensitivity for action modeling.
3.2 Stage 2: Aligning Video to Audio-Speech-Text
We exploit the correspondence between video and audio, speech, and text to align InternVideo2 to semantics explicitly. In structure, though InternVideo2 has a huge video encoder, its employed audio and text encoders are relatively lightweight. The used audio encoder is a 12-layer transformer initialized with BEATs Chen et al. (2023d) (90M). It takes in audio features, which are 64-dimensional log Mel filterbank spectrograms using a 25ms Hamming window, transformed from 10-second-long clips, padding with zeros. For the text and speech encoders, we initialize the text encoder and multimodal decoder using Bert-Large Devlin et al. (2018). Specifically, we utilize the initial 19 layers of Bert-Large as the text encoder, with the subsequent 5 layers equipped with cross-attention layers serving as the multimodal decoder.
For pretraining objectives, we establish alignment and fusion across modalities via text, including video, audio, image, and speech. We employ crossmodal contrastive and matching losses with masked language reconstruction one as:
where , , and are hyperparameters to tradeoff different losses. The employed and are standard and referenced from Cheng et al. (2022). Specifically, the crossmodal contrastive learning is given as:
where and denote the learned video and text embeddings, respectively. and indicates the modality of input signals and the text descriptions describing, respectively. computes the cosine similarity between two features. is the learnable temperature.
For the matching part, it is given as:
where computes the matching likelihood between video and text , respectively. denotes whether the given video and text are paired () or not ().
The employed masked language modeling loss is:
where computes the likelihood of the text token based on the previous ones. Here refers to video captions.
For training efficiency, we employ masking learning, aligning unmasked video tokens to tokens from other modalities first, then using full video tokens reconstruction shortly. Specifically, it consists of two steps as follows:
Aligning Masked Visual-Language-Audio.
We freeze the audio encoder, focusing primarily on aligning visual and text features. We employ the complete set of image, video, and audio-video data for pre-training, where in this training (the combinations denote the concatenated features from the used modalities).
Unmasked Visual-Audio-Language Post-Pretraining.
We freeze the vision encoder to achieve joint alignment of audio, visual, and text features. We conduct post-pretraining using a small subset of image and video data (25M), along with the complete set of audio (0.5M) and audio-video data (50M). Given that the parameters of the largest ViT-6B model are frozen, we abandon masking strategies here for the consistency with the inference, minimizing the degradations on downstream performance brought by masking. Here .
3.3 Stage3: Predicting Next Token with Video-Centric Inputs
To further enrich the semantics embedded in InternVideo2 and improve its user-friendly in human communications, we tune it by incorporating it into a VideoLLM with a LLM and a video BLIP Li et al. (2022b, c). We employ the progressive learning scheme in Li et al. (2023b) by using InternVideo2 as the video encoder and train a video blip for communicating with open-sourced LLM Zheng et al. (2023); Jiang et al. (2023). In training, the video encoder will be updated. Detailed training recipts are in Li et al. (2023b).
| Pretraining Stage | Dataset | Domain | # of clips | Annotation |
| Stage 1 | KMash | Web Video | 2M | Action Label |
| Stage 2 (img-txt) | LAION, etc | Web Image | 300M | Alt-text / Generated Caps |
| Stage 2 (vid-txt) | WebVid2M | Web Video | 250k | Alt-text |
| WebVid10M | Web Video | 9.7M | Alt-text | |
| InternVid | Youtube Video | 40M | Generated Caption | |
| Self-collected | Youtube Video | 60M | Generated Caption | |
| Stage 3 | LLaVA, etc | Web Image/Video | 2.1M | Conversation, QA |
4 Multimodal Video Data
We prepare our training data (video-only, video multimodal versions, and visual instructions) as given in Tab. 1.
4.1 Video-only Data for Masked AutoEncoders
We curate a video collection without labels and text named K-Mash from renowned action recognition datasets Carreira and Zisserman (2017); Goyal et al. (2017b); Monfort et al. (2020); Heilbron et al. (2015); Zhao et al. (2019), as detailed in Supp. It encompasses a wide range of video types, including first- and third-person perspectives, with both short and long durations, and featuring various characters and settings. Further, we give K-Mash with additionally sourced and meticulously selected 844K videos from YouTube for diversity. We observe notable action related task performance improvements in finetuned settings using K-Mash, as given in Sec. 5.4.2.
4.2 Videos with Audio-Speech Modalities for Contrastive Learning
We build a multimodal video dataset MVid with video-audio-speech information and their descriptions for pushing video perception via other modalities. It consists of 110M videos along with their VAS captions. In MVid, we collect videos from several sources (detailed in the supplementary material), segment them into clips, and automatically annotate them based on their solo unimodal or crossmodal inputs. We highlight the importance of temporal segmentation in clip generation and our video mutimodal annotation system. We find refining them leads to notable downstream improvements.
Temporal Consistency Matters.
We employ a temporal boundary detection model AutoShot Zhu et al. (2023c) to segment videos into clips instead of SceneDet filter from FFMPEG. It predicts boundaries based on temporal semantic variations rather than pixel differences, leading to generating semantically complete cuts without mixing extra frames with inconsistent context.
Video Multimodal Annotation.
We design a video multimodal annotation system VidCap to give proper unimodal and crossmodal descriptions for textualizing videos from different perceptions. It automatically captions visual, audio, and speech of MVid, then it corrects them and fuses them for cross-modal captions via LLM. The system frame is given in Fig. 3. VidCap has independent video, audio, speech captioner, and a LLM for caption refinement and fusion. For video, speech, and caption postprocessing, we employ existing methods as the video captioning pipeline in Wang et al. (2023d), WhisperV2-large model Radford et al. (2023), and Vicuna-1.5 Zheng et al. (2023). For audio, we craft a audio captioner upon VideoChat Li et al. (2023d), as we find no open-sourced one. It extracts audio features from inputs by Beats Chen et al. (2023d). We learn it by only tuning its Qformer using a combination of the large-scale audio-text corpus WavCaps Mei et al. (2023) dataset. Details are given in the supplementary material.
4.2.1 Video Instruction-Tuning Data for Next Token Prediction
We employ the training part of MVBench Li et al. (2023b). It comprises around 2M samples (both images and videos) from 34 distinct sources. This training dataset covers key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question anwser, 4) reasoning, and 5) classification.
5 Experiments
In our evaluation of InternVideo2, we assess the model performance of the three learning stages. The evaluation covers a wide range of tasks, including action recognition, video retrieval, question-answering, and more. It includes various scenarios such as zero-shot learning, finetuning, and others.
For InternVideo2 trained in stage 1, 2, and 3, we denote them with InternVideo2, InternVideo2, and InternVideo2, respectively. We also learn a CLIP-style InternVideo2 indicated by InternVideo2. It is post-pretrained from InternVideo2 by only preserving video and text encoders and contrastive loss.
Each training stage of InternVideo2-6B utilizes different configurations. In the first stage, we employ 256 NVIDIA A100 GPUs and train the model for 18 days. The second stage also utilizes 256 A100 GPUs and spans a training period of 14 days. Finally, in the third stage, we use 64 A100 GPUs and train the model for 3 days. We introduce DeepSpeed and FlashAttention Dao et al. (2022) for training and inference. Details and more results are given in the supp.
5.1 Video Classification
5.1.1 Action Recognition
We test InternVideo2 on Kinetics (i.e., K400, 600 and 700 Carreira and Zisserman (2017); Carreira et al. (2018, 2019)), Moments in Time V1 (MiT) Monfort et al. (2020), Something-Something V2 (SSv2) Goyal et al. (2017b), UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), Charades Gao et al. (2017), ActivityNet Heilbron et al. (2015) (ANet) and HACS Zhao et al. (2019). We evaluate in four settings: (a) end-to-end finetuning the whole backbone; (b) attentive probing is similar to linear pooling, but extra trains the attention pooling layer Yu et al. (2022). (c) linear probing which freezes the backbone and only trains the task head; and (d) zero-shot.
| Method | Training Data | Setting | K400 | K600 | K700 | SthSthv2 | MiT | ANet | HACS |
| CoVeR Zhang et al. (2021) | IV-3B | 16 448 | 87.1 | 87.9 | 79.8 | 70.8 | 46.1 | - | - |
| Hiera-H Ryali et al. (2023) | V-0.25M | 16 224 | 87.8 | 88.8 | 81.1 | 76.5 | - | - | - |
| CoCa-g Yu et al. (2022) | I-3B | 16 576 | 88.9 | 89.4 | 82.7 | - | 49.0 | - | - |
| MTV-H Wang et al. (2023a) | IV-370M | 32 280 | 89.9 | 90.3 | 83.4 | - | - | - | - |
| VideoMAEv2-g Wang et al. (2023a) | V-1.35M | 64 266 | 90.0 | 89.9 | - | 77.0 | - | - | - |
| V-JEPA-H Bardes et al. (2024) | V-2M | 16 224 | - | - | - | 77.0 | - | - | - |
| MVD-H Wang et al. (2023b) | IV-1.25M | 16 224 | - | - | - | 77.3 | - | - | - |
| UniFormerV2-L Li et al. (2022a) | IV-401M | 64 336 | 90.0 | 90.1 | 82.7 | 73.0 | 47.8 | 94.7 | 95.4 |
| InternVideo Wang et al. (2022) | V-12M | ensemble | 91.1 | 91.3 | 84.0 | 77.2 | - | - | - |
| InternVideo2-1B | IV-1.1M | 8 224 | 91.3 | 91.4 | 85.0 | 77.1 | 50.8 | - | - |
| InternVideo2-1B | IV-1.1M | 16 224 | 91.6 | 91.6 | 85.4 | 77.1 | 50.9 | - | - |
| InternVideo2-6B | IV-2M | 8 224 | 91.9 | 91.7 | 85.7 | 77.5 | 51.0 | - | - |
| InternVideo2-6B | IV-2M | 16 224 | 92.1 | 91.9 | 85.9 | 77.4 | 51.2 | 95.9 | 97.0 |
End-to-end Finetuning.
Tab. 2 shows InternVideo2-6B obtains new state-of-the-art (SOTA) results on Kinetics (92.1%/91.9%/85.9% on K400/600/700, respectively), SthSthv2, MiT, ANet, and HACS with only 16 frames, while the previous SOTAs require larger resolution (224 vs. 576) or model ensemble. As for MiT in Table 2, InternVideo2-6B exceeds the previous SOTA, CoCa-g, by a significant margin of 2.2% (51.2% vs. 49.0%). Regarding the temporal-related actions in Table 2, our InternVideo2-6B also surpasses MVD Wang et al. (2023b) on SSv2 (77.5% vs. 77.3%). Moreover, our InternVideo2-6B showcases top performance on untrimmed video analysis, as indicated in Table 2, with 95.9% on ActivityNet and 97.0% on HACS. These results affirm our model’s superior capability for robustly identifying complex actions across varied scenes.
| Method | Training Data | Setting | K400 | K600 | K700 | MiT | SSV2 |
| UMT-L | IV-25M | - | 82.8 | - | - | 40.3 | 54.5 |
| VideoMAEv2-g | V-1.35M | - | 82.1 | - | - | 35.0 | 56.1 |
| V-JEPA-H Bardes et al. (2024) | V-2M | 16 384 | 81.9 | - | - | - | 72.2 |
| DINOv2-g Oquab et al. (2023) | I-142M | 16 224 | 83.4 | - | - | - | 50.0 |
| VideoPrism-g Zhao et al. (2024) | V-619M | 16 288 | 87.2 | - | - | 45.5 | 68.5 |
| ViT-e Dehghani et al. (2023) | I-4B | 128 224 | 86.5 | - | - | 43.6 | - |
| ViT-22B Dehghani et al. (2023) | I-4B | 128 224 | 88.0 | - | - | 44.9 | - |
| CoCa-g Yu et al. (2022) | I-3B | 16 576 | 88.0 | 88.5 | 81.1 | 47.4 | - |
| InternVideo2-1B | IV-25.5M | 16 224 | 87.9 | 88.0 | 79.5 | 46.3 | 67.3 |
| InternVideo2-6B | IV-400M | 16 224 | 88.8 | 89.1 | 81.0 | 47.8 | 67.7 |
Attentive Probing.
As in Tab. 3, InternVideo2-6B not only outperforms ViT-22B Dehghani et al. (2023) and CoCa-g Yu et al. (2022) in scene-focused datasets but also surpasses or matches the performance of the latest video foundation model Bardes et al. (2024); Zhao et al. (2024), on datasets emphasizing temporal dynamics (SthSthV2). This underscores our model’s exceptional ability to understand and interpret both spatial and temporal information effectively.
| Method | Setting | K400 | SSV2 | UCF-101 | HMDB-51 |
| VideoMAEv2-H Wang et al. (2023a) | 12 224 | 25.8 | - | 56.4 | 34.1 |
| TVTSv2-H Zeng et al. (2023b) | 12 224 | 73.1 | - | 91.8 | 65.7 |
| OpenCLIP-G Cherti et al. (2023) | 8 224 | 78.3 | 35.8 | 90.7 | - |
| DINOv2-g Oquab et al. (2023) | 8 224 | 78.4 | 38.3 | 91.2 | - |
| InternVideo2-1B | 16 224 | 81.6 | 46.3 | 96.0 | 71.6 |
| InternVideo2-6B | 16 224 | 82.0 | 47.8 | 96.3 | 71.8 |
| InternVideo2-6B | 16 224 | 84.2 | 56.7 | 97.3 | 80.7 |
Linear Probing.
In Tab. 4, InternVideo2-1B significantly outperforms the previous SOTA, DINOv2-g Oquab et al. (2023), by notable margins: +3.2% on K400, +8.0% on SthSthV2, and +4.8% on UCF-101. As we scale the model, an upward trend in results is observed, underscoring the benefits of model enhancement. Notably, the integration of multimodal pretraining (stage 2) yields further rise in results. We suppose stage 2 enhances feature discrimination.
| Method | #F | Training Data | UCF | HMDB | MiT | SSv2-MC | Charades |
| CLIP Radford et al. (2021a) | 12 | I-400M | 68.9 | 43.2 | - | 29.6 | - |
| TVTSv2 Zeng et al. (2023b) | 12 | V-8.5M | 78.0 | 52.1 | - | 48.4 | - |
| VideoCoCa-g Yan et al. (2022) | 16 | V-145M | - | - | - | - | 25.8 |
| VideoPrism-g Zhao et al. (2024) | 16 | V-619M | - | - | - | - | 32.4 |
| InternVideo2-1B | 8 | IV-25.5M | 88.8 | 53.9 | 31.6 | 61.5 | 32.9 |
| InternVideo2-6B | 8 | IV-400M | 89.5 | 56.7 | 32.9 | 63.5 | 34.6 |
| Method | #F | K400 | K600 | K700 | |||
| top-1 | AVG | top-1 | AVG | top-1 | AVG | ||
| CLIP Radford et al. (2021a) | 8 | 58.4 | 70.1 | 55.1 | 67.2 | 46.1 | 58.4 |
| EVA-CLIP-L Sun et al. (2023c) | 1 | - | 65.0 | - | 64.9 | - | 59.1 |
| EVA-CLIP-E Sun et al. (2023c) | 1 | - | 69.8 | - | 69.3 | - | 63.4 |
| ViCLIP-L Wang et al. (2023d) | 8 | 64.8 | 75.7 | 62.2 | 73.5 | 54.3 | 66.4 |
| VideoCoCa-g Yan et al. (2022) | 16 | 72.0 | 81.3 | - | |||
| InternVL-6B Chen et al. (2023a) | 8 | 69.1 | 79.4 | 68.9 | 78.8 | 60.6 | 71.5 |
| EVA-CLIP-18B Sun et al. (2024) | 16 | 79.4 | - | 79.4 | - | 72.2 | |
| VideoPrism-g Zhao et al. (2024) | 16 | 76.4 | 85.4 | - | |||
| InternVideo2-1B | 8 | 73.1 | 82.4 | 72.8 | 81.8 | 64.9 | 75.2 |
| InternVideo2-6B | 8 | 72.7 | 82.2 | 71.7 | 81.2 | 64.2 | 75.2 |
Zero-shot.
Table 5 and 6 show InternVideo2 gets 72.7% / 71.7% / 64.2% on K400/600/700, respectively, outperforming others but VideoPrism on K400 (76.4%). On UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), MiT Monfort et al. (2020), SSv2-MC, and Charades, InternVideo2 gives an cutting edges over others. The clear gap between VideoPrism and InternVideo2 on K400 may signify the importance of pretraining corpus in VideoPrism (311M videos with text and 36.1M of them are manually labeled) for K400 in zero-shot. Note that on the datasets of Kinetics, UCF101 and HMDB51, which have a distribution closer to the pre-training dataset used in stage1, the performance of Internvideo2-6B is slightly inferior to that of Internvideo2-1B. We suppose this is caused by Internvideo2-6B uses a more abundant pretraining dataset in stage2, leading to the forgetting of pretraining data in stage1.
5.1.2 Temporal Action Localization
We evaluate models on four temporal action localization (TAL) datasets: THUMOS14 Idrees et al. (2017), ActivityNet Krishna et al. (2017), HACS Segment Zhao et al. (2019) and FineAction Liu et al. (2022) in a feature-based manner with finetuning. We employ output of the 7-th layer from InternVideo2 for inputs as the corresponding features without fusing anything else. ActionFormer Zhang et al. (2022) is used as the detection head. We report mean Average Precision(mAP) under multiple tIoU as in Lin et al. (2019); Chen et al. (2022b); Zhang et al. (2022); Yang et al. (2023). In Table 7, InternVideo2-6B gets the highest mAP among all comparisons in all datasets, while InternVideo2-1B nearly surpass other methods except in THUMOS14. We find InternVideo2-6B almost consistently improves mAP with a notable margin from InternVideo2-1B except in FineAction. We suppose scaling model capacity without data refinement cannot nontrivially improve fine-grained discrimination abilities of models. Scaling detailed annotations in training may address this issue.
5.1.3 Video Instance Segmentation
We evaluate on the popular Video Instance Segmentation (VIS) dataset, Youtube-VIS 2019 Yang et al. (2019). Built upon Mask2Former Cheng et al. (2021), we employ the video encoder of InternVideo2 as backbone with ViT-adapter Chen et al. (2022c) for features. We also try InternViT Chen et al. (2023a) for comparisons. In Table 8, InternVideo2 gets the highest mAP among all. It improves mAP with a nontrivial margin over InternViT and Swin-L. This validates its effectiveness in relatively fine-grained spatiotemporal perception.
5.2 Video-Language Tasks
We evaluate InternVideo2 on video retrieval, captioning, and multi-choice question-answersing (QA). The former two tasks are conducted by matching video representation and the candidate text ones using the text encoder in stage 2. The latter is tested by the VideoLLM learned in stage 3. We also test audio tasks.
| Method | MSR-VTT | LSMDC | DiDeMo | MSVD | ANet | VATEX | ||||||
| T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | |
| CLIP Radford et al. (2021a) | 30.4 | 24.2 | 13.9 | 11.9 | 12.7 | 18.7 | 40.5 | 57.2 | 9.1 | 13.2 | - | - |
| CLIP4Clip Luo et al. (2022) | 32.0 | - | 15.1 | - | - | - | 38.5 | - | - | - | - | - |
| ViCLIP Wang et al. (2023d) | 42.4 | 41.3 | 20.1 | 16.9 | 18.4 | 27.9 | 49.1 | 75.1 | 15.1 | 24.0 | - | - |
| InternVideo-L Wang et al. (2022) | 40.7 | 39.6 | 17.6 | 13.2 | 31.5 | 33.5 | 43.4 | 67.6 | 30.7 | 31.4 | 49.5 | 69.5 |
| UMT-L Li et al. (2023a) | 40.7 | 37.1 | 24.9 | 21.9 | 48.6 | 49.9 | 49.0 | 74.5 | 41.9 | 39.4 | - | - |
| VideoCoCa-g Yan et al. (2022) | 34.4 | 64.7 | - | - | - | - | - | - | 34.5 | 33.0 | 53.2 | 73.6 |
| VideoPrism-g Zhao et al. (2024) | 39.7 | 71.0 | - | - | - | - | - | - | 52.7 | 50.3 | 62.5 | 77.1 |
| InternVideo2-1B | 51.9 | 50.9 | 32.0 | 27.3 | 57.0 | 54.3 | 58.1 | 83.3 | 60.4 | 54.8 | 70.4 | 85.4 |
| InternVideo2-6B | 55.9 | 53.7 | 33.8 | 30.1 | 57.9 | 57.1 | 59.3 | 83.1 | 63.2 | 56.5 | 71.5 | 85.3 |
| Method | MSR-VTT | LSMDC | DiDeMo | MSVD | ANet | VATEX | ||||||
| T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | T2V | V2T | |
| CLIP Radford et al. (2021a) | 38.2 | 38.7 | 22.5 | 22.6 | 32.2 | 33.9 | - | - | 26.1 | 26.9 | - | - |
| CLIP4Clip Luo et al. (2022) | 45.6 | 45.9 | 24.3 | 23.8 | 43.0 | 43.6 | 45.2 | 48.4 | 40.3 | 41.6 | - | - |
| ViCLIP Wang et al. (2023d) | 52.5 | 51.8 | 33.0 | 32.5 | 49.4 | 50.2 | - | - | 49.8 | 48.1 | - | - |
| UMT-L Li et al. (2023a) | 58.8 | 58.6 | 43.0 | 41.4 | 70.4 | 65.7 | 58.2 | 82.4 | 66.8 | 64.4 | 72.0 | 86.0 |
| InternVideo2-6B | 62.8 | 60.2 | 46.4 | 46.7 | 74.2 | 71.9 | 61.4 | 85.2 | 74.1 | 69.7 | 75.5 | 89.3 |
5.2.1 Video Retrieval
We evaluate the video retrieval on six popular benchmarks Heilbron et al. (2015); Xu et al. (2016); Rohrbach et al. (2015); Anne Hendricks et al. (2017); Chen and Dolan (2011), as shown in Tab. 9 and 10. In eval, eight frames from the input videos are uniformly sampled. We report R@1 scores for both text-to-video (t2v) and video-to-text (v2t) tasks in Tab. 9 and 10. R@5 and R@10 are given in Supp.
Table 9 and 10 demonstrate that InternVideo2 outperforms other state-of-the-arts with a notable margin in both t2v and v2t of all used datasets no matter in zero-shot or finetuned settings, except for the v2t of MSR-VTT, where VideoPrism gives the best result. This shows the video-language semantic alignment of transferrity of InternVideo2.
| Method | MSR-VTT | VATEX | MSVD | ANet Cap |
| BLIP2 Li et al. (2022b) | 19.0 | - | 27.4 | - |
| Video-LLaMA Zhang et al. (2023a) | 29.1 | - | 47.0 | - |
| Student Captioning Model Chen et al. (2024b) | 31.5 | - | 49.2 | - |
| VideoCoCa Yan et al. (2022) | 27.1 | 22.8 | - | 19.3 |
| Flamingo-3B Alayrac et al. (2022) | - | 40.1 | - | - |
| Flamingo-9B Alayrac et al. (2022) | - | 39.5 | - | - |
| VideoPrism-B Zhao et al. (2024) w/ PaLM-1B | 40.3 | 24.2 | - | - |
| VideoPrism-B Zhao et al. (2024) w/ PaLM-8B | 38.5 | 31.7 | - | - |
| InternVideo2-1B | 43.5 | 49.2 | 93.1 | 19.6 |
5.2.2 Video Captioning
As shown in Tab. 11,we evaluate zero-shot video captioning using InternVideo2 learned in stage 3 on MSR-VTT, VATEX, MSVD, and ANet Captions. InternVideo2 delivers superior CIDEr scores than VideoCoCa, Flamingo, and VideoPrism. Despite models scale, it may suggest a progressive learning scheme involving masked reconstruction, crossmodal contrastive learning, and next token prediction would present better video-to-text generation performance than those which only utilize some of them.
5.2.3 Multi-Choice Video Question Answering
We evaluate zero-shot multi-choice (MC) video QA using InternVideo2 in stage 2 on MSR-VTT, VATEX, and EgoSchema Mangalam et al. (2023). Note EgoSchema is a long-form egocentric video benchmark in the form of video QA. Tab. 12 shows InternVideo2 consistently improves MC accuracy compared with previous SOTAs except on LSMDC, where it gets a comparable result with InternVideo. Its zero-shot result on EgoSchema increases previous state-of-the-art by 9 points, suggesting its better generalization on long video understanding than others.
5.2.4 Video Temporal Grounding
We evaluate InternVideo2 on two temporal grounding (VTG) datasets: Qvhighlight Lei et al. (2021) and Charade-STA Gao et al. (2017). The eval setting and used features are the same as in TAL. We use CG-DETR Moon et al. (2023) as the grounding head. We report R1@0.3, R1@0.5, R1@0.7, and mAP for moment retrieval as in Lei et al. (2021); Moon et al. (2023); Lin et al. (2023). Highlight Detection is evaluated in terms of “Very Good” mAP and HiT@1. In Table 13, InternVideo2-1B and InternVideo2-6B bring gradual performance improvements compared to CLIP Radford et al. (2021b) and CLIP Radford et al. (2021b)+Slowfast Feichtenhofer et al. (2019). This suggests that a larger spatiotemporal model is more beneficial to short-term video semantic alignment capabilities.
5.2.5 Audio-related Tasks
We evaluate InternVideo2’s audio and text encoders on audio tasks, including audio-text retrieval on AudioCaps Kim et al. (2019), Clothov1, and Clothov2 Drossos et al. (2020); audioQA on ClothoAQA Lipping et al. (2022) and Audio-MusicAVQA Behera et al. (2023); and audio classification on the ESC-50 Piczak (2015). As shown in Tab. 14, 15(a), and 15(b), our model achieves state-of-the-art performance on all downstream tasks. Considering the limited size of the used audio and text encoders, these audio-related results show crossmodal contrastive learning’s benefits are mutual to the used modalities. Audio and the corresponding text models also gain from this learning.
| Model | ViEncoder | LLM | MVBench | VideoChatGPT | MoVQA |
| GPT-4V OpenAI (2023b) | - | GPT-4 | 43.5 | - | - |
| Gemini Pro Team et al. (2023) | - | - | 37.7 | - | - |
| LLaVA Liu et al. (2023) | CLIP-L | Vicuna-7B | 36.0 | - | - |
| mPLUG-Owl-V Ye et al. (2023a) | CLIP-L | LLaMA-7B | 29.7 | - | 24.8 |
| VideoChatGPT Maaz et al. (2023a) | CLIP-L | Vicuna-7B | 32.7 | 2.38 | 22.9 |
| VideoLLaMA Zhang et al. (2023c) | EVA-CLIP-g | Vicuna-7B | 34.1 | 1.98 | - |
| VideoChat2 | UMT-L | Vicuna-7B | 51.1 | 2.98 | 34.7 |
| VideoChat2 | InternVideo2-1B | Mistral-7B | 60.9 | 2.67 | 41.0 |
| Model | Zero-shot | Finetuning | |
| Action Recognition | Video Retrieval | Action Recognition | |
| InternVideo2-1B | 55.5 | 55.0 | 73.2* |
| InternVideo2-6B | 56.9 | 56.9 | 73.6* |
| Model | Teacher | Data | K400 | K600 | K700 | MiT | SSv2 | Avg |
| ViT-L | CLIP-L | K710 | 90.3 | 90.4 | 83.2 | 48.0 | 74.7 | 77.3 |
| ViT-L | CLIP-L | K-Mash | 90.5 | 90.4 | 83.4 | 48.1 | 74.7 | 77.4 |
| ViT-1B | InternVL-6B | K710 | 90.9 | 91.0 | 84.7 | 49.8 | 75.9 | 78.5 |
| ViT-1B | InternVL-6B | K-Mash | 91.4 | 91.5 | 85.1 | 50.5 | 76.5 | 79.0 |
| ViT-1B | InternVL-6BVideoMAE-g | K-Mash | 91.3 | 91.4 | 85.0 | 50.8 | 77.1 | 79.1 |
| ViT-1B | InternVL-6B | K-Mash | 91.3 | 91.5 | 85.1 | 50.6 | 76.6 | 79.0 |
| ViT-6B | InternVL-6BVideoMAE-g | K-Mash | 91.9 | 91.7 | 85.7 | 51.0 | 77.5 | 79.6 |
5.3 Video-centric Dialogue and its Applications
Tab. 16 shows results on MVBench Li et al. (2023b), VideoChatGPT-Eval Maaz et al. (2023a), and MoVQA Zhang et al. (2023b) from VideoChat2 (a video-based large language model) Li et al. (2023b) with different backbones, compared with other MLLMs. Note VideoChat-InternVideo2 outperforms other systems with a notable margin, not only on the average score but also in each subtask (in the Supp), except in VideoChatGPT-Eval. Considering these benchmarks not only involve perception but also reasoning, these results suggest InternVideo2 does embed knowledge to partially model the world, at least compared with other models. It also validates the importance of learning a transferrable video representation for current video-related MLLM. We also give several qualitative evaluations with popular GPT-4V and GeminiPro, including dialogues on action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9).
5.4 Ablation Studies
5.4.1 Scaling Video Encoder
Tab. 17 gives InternVideo2’s average performance on action recognition and video retrieval. It shows that scaling video encoder from 1B to 6B still leads to notable improvements in generalization of action recognition and video-retrieval by 1.4% and 1.9% (in zero-shot), respectively. Meanwhile, the increase in finetuned action recognition result is relatively marginal (0.4%) with the growth of model scale.
5.4.2 Training Data and used Teachers in Stage 1
In Tab. 18, we examine the impact of distillation teachers and used dataset sizes on model performance. (a) Data Scale: Note that pretraining data scale for MAE should grow with the increasing model scale, otherwise the downstream performance would be saturated, such as K710 (0.66M videos) for ViT-L, K-Mash for ViT-1B, and K-Mash for ViT-6B. (b) Teacher: Tab. 18 reveals that the synergy between a multimodal teacher (e.g., CLIP Radford et al. (2021b)) and a motion-aware teacher (e.g., MAE Tong et al. (2022)) markedly boosts performance, especially on SthSthV2. It highlights the importance of strategic teacher model selection in the distillation process.
| Method | MSR-VTT |
| Baseline (video & text encoders w/ video-text learning) | 24.7 |
| Baseline + audio encoder w/ audio-text learning | 24.0 |
| Baseline + speech encoder w/ video-speech-text learning | 24.9 |
| Baseline + audio encoder w/ video-audio-text learning | 27.8 |
| Baseline + audio & speech encoders + video-audio-speech-text learning | 25.7 |
| SceneDet | AutoShot | Video Cap | Audio Cap | Speech Cap | MSR-VTT |
| ✓ | ✓ | 24.7 | |||
| ✓ | ✓ | ✓ | 26.6 | ||
| ✓ | ✓ | ✓ | ✓ | 27.1 | |
| ✓ | ✓ | ✓ | ✓ | 34.8 |
5.4.3 Training Arch, Method, and Data in Stage 2
We ablate the necessity of introducing an audio encoder in Stage 2. We employ ViT-B and Bert-B for video and text encoders, respectively. The used text are simple video captions. The baseline is conducting video-text contrastive and matching as well as masked language generation loss for training with only video and text encoders. Other settings including adding audio or speech encoder or them both, and how to update newly added encoders i.e., whether train them with only text encoder or both video and text encoders. Tab. 19 shows that only introducing audio encoder and learn it along with both video and text encoders can best improve video-text retrieval performance. The speech encoder harms such effectiveness more or less.
Further, we verify the impact of video temporal segmentation and the used captions as text inputs in Stage 2 in Tab. 20. We still use ViT-B and Bert-B for video-text training. Tab. 20 finds the fused text from video-audio-speech works best for retrieval tasks compard with others, rising zero-shot t2v R1 of MSR-VTT from 24.7 to 27.1. Moreover, using AutoShot instead of SceneDet notably improves t2v retrieval (increasing by nearly 7 points). It validates the effectiveness of the introduced video-text dataset and its annotation systems.
6 Conclusion and Discussion
We introduce a novel video foundation model called InternVideo2, which achieves state-of-the-art performance across various video and audio tasks. In InternVideo2, we combine masked video token reconstruction, video-audio-text contrastive learning, and next token prediction into a unified framework. Additionally, we create a new video-text dataset that incorporates video-audio-speech fused captions as descriptions. The dataset contains temporally segmented clips with semantic differences. These designs in InternVideo2 contribute to enhancing video understanding in both perception and reasoning tasks. Notably, InternVideo2 excels in video-related dialogue and long video understanding, demonstrating its effectiveness in various video-related tasks.
Limitations and Discussions.
Despite its achievements, InternVideo2 does not introduce novel training methods or architectural advances. Instead, it leverages existing learning techniques for scheme exploration while focusing on improving data processing to enhance spatiotemporal perception, semantic alignment, and basic knowledge embedding. Similarly to previous studies Li et al. (2023a); Ye et al. (2023b), the model still grapples with limitations stemming from fixed input resolutions, sampling rates, and highly compressed tokens, which restrict its ability to express rich video information and capture fine-grained details.
The progressive learning scheme adopted by InternVideo2 strikes a balance between model capabilities and training compute. While jointly learning the three optimization objectives simultaneously is computationally feasible, scalability becomes an issue when confronted with limited resources.
Although InternVideo2 has demonstrated leading performance in long video understanding and reasoning benchmarks, it cannot guarantee an implicit world model that ensures consistency in visual reasoning. The inherent constraints imposed by fixed input representations, coupled with the complexity of visual reasoning tasks, present challenges in achieving a comprehensive and consistent understanding of the visual world.
7 Broader Impact
It is important to acknowledge that, similar to other foundational models, InternVideo2 has the potential to embed biases present in its training data and the associated models used during training, such as neural teachers Tong et al. (2022); Chen et al. (2023a) and language models (LLMs) Jiang et al. (2023); Zheng et al. (2023). These biases may emerge due to a variety of factors, including the personal ideas, preferences, values, and perspectives of the data creators and the training corpus utilized.
The presence of biases in AI models can have societal implications and reinforce existing inequalities or prejudices. Biases within InternVideo2 could manifest in the form of unfair or discriminatory outputs, potentially perpetuating social biases or stereotypes present in the training data. Consequently, it is crucial to carefully consider the potential impact of deploying InternVideo2 in real-world applications and take proactive measures to mitigate biases and ensure fairness.
Appendix A Model
| Stage | ViT-6B | Output Size |
| Video | sparse sampling | 38224224 |
| Patch | 11414, 3200 | 32008256 |
| Embedding | stride 11414 | |
| Position | learnable, 3D sine-cosine initialization | 32002048 |
| Embedding | 32002048 | |
| Mask | semantic mask w/ mask ratio | 32002048(1-) |
| Encoder | 48 AttnPool(768) | 32002048(1-) 7681 |
| Projection | , , | 32002048(1-) 14082048(1-) 7681 |
Video Encoder.
In Tab. 21, we take ViT-6B as an example and omit the class token for a simple presentation. “MHSA”, “MLP”, “AttnPool”, and “LN” refer to spatiotemporal multi-head self-attention, multi-layer perception, attention pooling Yu et al. [2022] and root mean square layer normalization Zhang and Sennrich [2019]. and means the layer number for unmasked token alignment with multimodal and motion-aware teachers. We mark the channel number, frame number, spatial size, and token number by different colors. The projection layers are dropped after stage 1 training.
Appendix B Video-centric Multimodal Data
We prepare our training data according to the learning objectives of the three stages in our learning scheme. Specifically, it consists of video-only pretraining set for masked video token reconstruction, Video-Audio-Speech-Text one for multimodal alignment, and video instruction dataset for human-computer interaction alignment. They are detailed in the following.
B.1 Video-only Data
| Dataset | K710 | SthSthV2 | HACS | ANet | MiT | Self-collected |
| K-Mash | 658K | 169K | 106K | 15K | 152K | 0 |
| K-Mash | 658K | 169K | 106K | 15K | 207K | 844K |
To create the curated collection of videos, named K-Mash, we source videos from renowned action recognition datasets such as Kinetics-400 (K400) Carreira and Zisserman [2017], Something-Something (Sth) Goyal et al. [2017b], Moments in Time (MIT) Monfort et al. [2020], ActivityNet Heilbron et al. [2015], and HACS Zhao et al. [2019]. These datasets provide a wide range of video types, including both first-person and third-person perspectives, short and long durations, and featuring a rich variety of characters and settings.
For the enhanced version of the dataset, called K-Mash, we push a step further and meticulously selected an additional 844,000 videos from YouTube to further enhance the diversity of the dataset. It’s important to note that all videos in this dataset are utilized for training without any labels, allowing the model to learn from the unlabeled data in an unsupervised manner. This approach helps to broaden the model’s understanding of different visual concepts and improves its performance on various video-related tasks.
B.2 Videos with Audio-Video-Speech Modalities
In addtion to publicly available video-text datasets (e.g. InternVid [Wang et al., 2023d] and WebVid [Bain et al., 2021]), we introduce a new video dataset that incorporates both audio-visual-speech information and their corresponding textual descriptions. This dataset is included in the training process of InternVideo2. To create this multimodal dataset, named MVid, we leverage several video sources and provide detailed annotations. MVid includes videos with synchronized audio, visual, and speech information, along with their corresponding textual descriptions. This multimodal dataset enables the training of InternVideo2 to better understand and capture the connections between different modalities, enhancing its performance in various video-related tasks that require audio, visual, and speech understanding.
Collection. In the MVid dataset, approximately half of the videos are sourced from YouTube, while the remaining videos are gathered from anonymous sources. This is to improve the diversity of dataset, as relying solely on YouTube may result in limited depth of the dataset. Furthermore, to study the impact of video culture backgrounds on the learned models, a small portion of the dataset consists of Chinese data. These videos were collected with proper permissions for academic usage, ensuring compliance with legal and ethical considerations.
By incorporating videos from various sources and including a subset of Chinese data, MVid provides a more diverse and representative dataset for training InternVideo2. This approach allows the model to learn from a wide range of video content, encompassing different cultural backgrounds and further enhancing its ability to understand and process videos from various sources.
Trimming. In our approach, instead of relying on the widely-used SceneDet filter of FFMPEG, we employ a temporal boundary detection model called AutoShot Zhu et al. [2023c] to segment videos into clips. AutoShot is capable of predicting clip boundaries based on temporal semantic variations, as opposed to pixel differences. This leads to the generation of semantically complete cuts without mixing extra frames that may contain inconsistent context. By using AutoShot, we aim to reduce captioning errors by producing fewer clips with obvious transitions, resulting in a more coherent input for video captioning models. In the inference of AutoShot, we use a threshold of 0.5 to determine the shot boundaries for AutoShot’s estimations.
For the video dataset, we first preserve clips longer than 2 seconds. For video clips longer than 30 seconds, as the segments within the clip are from the same shot, we randomly choose a 30-second segment. During this process, we also abandon clips with still or extreme dynamics, such as browsing a photo gallery.
Annotation. We automatically caption visual, audio, and speech of MVid. Then we correct them and fuse them for cross-modal captions for training using LLM. We list several annotation examples of our method in Fig. 10.
Audio Captioner. We craft a audio captioner upon VideoChat [Li et al., 2023d], as we find no reliable ones. It extracts audio features from inputs by Beats [Chen et al., 2023d]. We learn it by only tuning its Qformer (the interface between audio encoder and LLM) using a combination of the large-scale audio-text corpus WavCaps [Mei et al., 2023] dataset.
Speech Captioner. We utilize the audio transcription model Whisper [Radford et al., 2023] to obtain speech from videos.Concretely, we use the WhisperV2-large model since its concurrent state-of-the-art performance. During the data collection process, a portion of the data is directly adopted from YT-Temporal-180M, which already has well-aligned timestamps and adjusted speech recognition content. The remaining data is first passed through a pre-trained language identification model Fasttext-lid [Joulin et al., 2016] to determine the language category, and then transcribed the non-English text into English using the pretrained Seamless M4T [Communication et al., 2023] model. For the text with language identification confidence less than 0.95, we use Vicuna-1.5 as a translation alternative.
Caption Triming & Fusion with LLM. After obtaining the captions of audio, video, and speech from the given video, we utilize an LLM (Vicuna-1.5 [Zheng et al., 2023]) to integrate the uni-modal captions into the multimodal ones. To fulfill the request for multiple contrastive objectives, we combine the audio caption with the video caption as the audio-visual caption, as well as integrate the audio, video, and speech captions as the audio-visual-subtitle captions. In this way, we acquire 5 types of captions (3 uni-modal captions (A, V, S) and 2 multi-modal captions (AV, AVS)) for each video automatically. Specifically, we have carefully designed prompt templates (Fig. 23) and employed vLLM [Kwon et al., 2023] for inference acceleration, effectively get the visual caption, audio caption, subtitle, audio-visual caption and audio-visual-speech caption while maintaining a natural human-like subtitle style.
Filtering & Sampling. After obtaining the captions, we calculate the CLIP similarity between the video segments and captions. We select the top 60 million data as the video segment data for MVid. For LAION-2B, we only select samples with CLIP similarity in the top 158 million for training.
B.3 Instruction-Tuning Data for Video Dialogues
We employ the training part of MVBench [Li et al., 2023b]. It comprises 1.9M samples (both images and videos) from 34 distinct sources. They are sampled from the public image / video datasets and reorganized in a standard question-anwser format. This training dataset covers key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question anwser, 4) reasoning, and 5) classification.
Appendix C Experiments
C.1 Ablations
C.1.1 How InternVideo2 Works in Feature-based Tasks.
We study which part of InternVideo2’s predictions are suitable for feature-based tasks, i.e. temporal action localition. We adhere the same train and test protocols as in the main paper.
In Tab. C.1.1, the most effective feature tends to be located within the last few layers of the video encoder. This observation aligns with the similarity between feature-based temporal tasks and linear probing classification, which is reasonable. We undertake comprehensive experiments to investigate the impact of features from various layers, as detailed in Table 24. The results reveal the best features appear between the last 5-th layer and 7-th layer.
| Layer Index | THUMOS-14 | ActivityNet | HACS Segment | FineAction | ||||
| 1B@mAP | 6B@mAP | 1B@mAP | 6B@mAP | 1B@mAP | 6B@mAP | 1B@mAP | 6B@mAP | |
| -1 | 67.9 | 70.3 | 39.0 | 40.7 | 39.5 | 42.1 | 25.4 | 25.3 |
| -2 | 68.4 | 71.0 | 39.3 | 40.5 | 41.0 | 42.7 | 26.2 | 26.4 |
| -3 | 69.0 | 71.3 | 39.7 | 40.5 | 41.2 | 42.7 | 27.0 | 26.6 |
| -4 | 69.3 | 71.4 | 39.6 | 41.1 | 41.3 | 43.1 | 27.1 | 27.1 |
| -5 | 69.9 | 71.8 | 39.7 | 40.9 | 41.4 | 43.1 | 27.2 | 27.7 |
| -6 | 69.6 | 72.0 | 39.6 | 41.2 | 40.7 | 43.3 | 27.0 | 27.7 |
| -7 | 69.5 | 71.9 | 40.0 | 41.1 | 40.6 | 42.8 | 26.9 | 27.7 |
C.2 Video Retrieval
We detail R@1, R@5, and R@10 of zero-shot video retrieval from InternVideo2 in Tab. 25(c) -25(f) for reference.
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 51.9 | 74.6 | 81.7 | 49.6 | 73.6 | 81.2 |
| InternVideo2-1B | 8 | 51.9 | 75.3 | 82.5 | 50.9 | 73.4 | 81.8 |
| InternVideo2-6B | 4 | 54.5 | 77.5 | 83.7 | 52.3 | 75.3 | 83.5 |
| InternVideo2-6B | 8 | 55.9 | 78.3 | 85.1 | 53.7 | 77.5 | 84.1 |
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 31.5 | 51.3 | 59.5 | 27.1 | 44.8 | 51.8 |
| InternVideo2-1B | 8 | 32.0 | 52.4 | 59.4 | 27.3 | 44.2 | 51.6 |
| InternVideo2-6B | 4 | 34.8 | 54.0 | 61.6 | 30.1 | 48.0 | 55.0 |
| InternVideo2-6B | 8 | 33.8 | 55.9 | 62.2 | 30.1 | 47.7 | 54.8 |
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 70.7 | 93.7 | 96.9 | 85.9 | 97.6 | 99.2 |
| InternVideo2-1B | 8 | 70.4 | 93.4 | 96.9 | 85.4 | 97.6 | 99.1 |
| InternVideo2-6B | 4 | 71.1 | 93.8 | 97.0 | 85.2 | 97.7 | 99.4 |
| InternVideo2-6B | 8 | 71.5 | 94.0 | 97.1 | 85.3 | 97.9 | 99.3 |
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 56.7 | 78.7 | 83.9 | 54.4 | 74.4 | 80.6 |
| InternVideo2-1B | 8 | 57.0 | 80.0 | 85.1 | 54.3 | 77.2 | 83.5 |
| InternVideo2-6B | 4 | 56.2 | 77.6 | 83.6 | 53.2 | 76.8 | 82.7 |
| InternVideo2-6B | 8 | 57.9 | 80.0 | 84.6 | 57.1 | 79.9 | 85.0 |
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 56.9 | 81.7 | 89.8 | 53.6 | 80.0 | 88.5 |
| InternVideo2-1B | 8 | 60.4 | 83.9 | 90.8 | 54.8 | 81.5 | 89.5 |
| InternVideo2-6B | 4 | 59.4 | 83.2 | 90.3 | 53.7 | 80.5 | 88.9 |
| InternVideo2-6B | 8 | 63.2 | 85.6 | 92.5 | 56.5 | 82.8 | 90.3 |
| Method | #F | Text-to-Video | Video-to-Text | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| InternVideo2-1B | 4 | 58.9 | 83.0 | 88.7 | 83.6 | 94.8 | 97.0 |
| InternVideo2-1B | 8 | 58.1 | 83.0 | 88.4 | 83.3 | 94.3 | 96.9 |
| InternVideo2-6B | 4 | 59.8 | 84.2 | 89.7 | 82.5 | 94.6 | 97.2 |
| InternVideo2-6B | 8 | 59.3 | 84.4 | 89.6 | 83.1 | 94.2 | 97.0 |
| Method | Backbone | Synopsis | Temporal | Spatial | Causal | Hypothetical | Knowledge | Single-Scene | Multi-Scene | Full-Scene | Overall |
| Mplug-Owl Ye et al. [2023b] | CLIP ViT-L | 25.1 | 19.9 | 25.3 | 21.9 | 23.5 | 27.5 | 25.2 | 23.5 | 22.1 | 24.8 |
| Otter Li et al. [2023f] | CLIP ViT-L | 22.6 | 20.7 | 19.6 | 26.1 | 24.2 | 21.8 | 23.1 | 22.1 | 21.3 | 22.6 |
| VideoChatGPT Maaz et al. [2023b] | CLIP ViT-L | 23.8 | 20.2 | 22.1 | 22.1 | 21.4 | 24.1 | 23.4 | 22.7 | 22.3 | 22.9 |
| VideoChat Li et al. [2023d] | Eva-g | 33.6 | 24.3 | 34.5 | 36.6 | 35.5 | 32.0 | 35.3 | 32.9 | 33.3 | 34.7 |
| VideoChat2 | InternVideo2 | 42.6 | 27.8 | 39.9 | 44.3 | 442.5 | 41.2 | 40.9 | 39.3 | 38.6 | 40.1 |
C.3 Movie Understanding.
We evaluate InternVideo2 on the MoVQA for movie understanding. It is a long-form movie question-answering datasetZhang et al. [2023b]. MoVQA assesses the diverse cognitive capabilities of multimodal systems by considering both video length and clue length, relying on multi-level temporal lengths (single-scene, multi-scene and full-scene). There are six types of QAs, including information synopsis, temporal perception, spatial perception, causal reasoning, hypothetical reasoning and external knowledge. We evaluate InternVideo2-6B in the form of open-ended QAs, and detailed results are shown on Table 26.
C.4 Video-Centric Dialogue System
We conduct qualitative evaluations of InternVideo2 on several challenging tasks (most videos are from Perception Test Patraucean et al. [2024]) such as action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9). Compared to other models, our model provides more concise and accurate answers.